# Machine Learning with Physics Knowledge for Prediction: A Survey

### v1

@@ 4.2 Meta Learning
Whereas multi-task learning aims to exploit common knowledge while learning a given set of tasks in parallel, meta learning aims to optimize a learning procedure on the given tasks that can later be applied to novel tasks. Both paradigms can leverage a set of source tasks to learn better solutions on target tasks and are therefore closely related to transfer learning Zhuang et al. (2021b). However, whereas multi-task learning optimizes source and target tasks jointly, meta learning considers the source tasks only during pre-training allowing for computationally efficient learning on the target tasks. In general, the result of meta-learning could be a complete algorithm. However, the methods discussed in the following are content with learning suitable initial parameters or other hyperparameters for given algorithms.
@@
Several works have applied these techniques for learning dynamics models (Lin et al., 2020; Qin et al., 2022). Meta-L (Lin et al., 2020) uses MAML for linear time-variant (LTV) systems. The model is optimized to achieve small prediction error after updating the linear parameters of the dynamics with a single gradient step. During deployment, the fast adaptation makes it possible to adapt to changes of the dynamics online. MetaPDE (Qin et al., 2022) employs a similar approach to nonlinear dynamics. They compare MAML with a slightly different meta learning technique, LEAP (Flennerhag et al., 2019), for meta learning a PINN that can be quickly adapted to different PDE parameters. While LEAP was faster during meta-learning, MAML performed faster in deployment. NRPINN (Liu et al., 2022) is a similar approach that uses REPTILE (Nichol et al., 2018) on a modified PINN loss, which contains an additional term that can make use of labeled data, if available.
@@
Another MAML-based approach is used by iMODE (Li et al., 2023), which adapts a latent input to a more general PINN that is shared among different PDE instances. A bi-level optimization problem is solved to optimize both, the common PINN network and the initialization of its latent input. Furthermore, PCA is applied to learn a compressed form of the latent input. When examples with labeled PDE parameters are available during training, iMode can learn a diffeomorphism between the labeled parameters and the compressed latent inputs, which can be used for system identification. DyAd (Wang et al., 2022d) directly feeds the parameters of the PDE as input to the model. The architecture is geared towards 2D images/velocity fields. It uses 3D convolutions and AdaPad, which predicts boundary conditions using a nonlinear transformation of the latent used for padding the images. As the PDE parameters are often unknown, DyAd also trains an encoder network that predicts the parameters from a history of states. The encoder is trained using weak supervision, e.g. by exploiting when a subset of the parameters are known.
@@
Instead of predicting latent inputs, hypernetworks can be used to predict the weights of the common PINN. For example, HyperPINN (de Avila Belbute-Peres et al., 2021) trains a hypernetwork that takes the parameters of a given class of PDEs (e.g. Burger’s equation) as input and outputs the parameters of a PINN. To reduce the number of the parameters of the PINN, Hyper-LR-PINNs (Cho et al., 2023b) restrict the weight matrices of the neural network to be low-rank, parameterized by the matrices of an SVD decomposition, that is, the weight matrix of layer $i$ is given by $\mathbf{W}_{i}=\mathbf{U}_{i}\bm{\Sigma}_{i}\mathbf{V}_{i}^{\top}$, with rectangular matrices $\mathbf{U}_{i}$ and $\mathbf{V}_{i}$, and a diagonal matrix $\bm{\Sigma}_{i}$. During offline pre-training, $\mathbf{U}$ and $\mathbf{V}$ are optimized directly along with the hypernetwork that predicts the elements of $\bm{\Sigma}$. During deployment, only the diagonal elements are optimized, using the hypernetwork for predicting their initial values.
@@
While the aforementioned methods were targeted towards a specific type of PDEs, some approaches aim to learn even more general networks. LEADS (Yin et al., 2021) learns a dynamics model, where the dynamics may also change over time. The model predicts the state derivatives and decomposes them into shared and environment-specific components. It is trained by minimizing the squared error between predicted derivatives and targets, with an additional convex regularizer penalizing the complexity of the environment-specific terms. Different regularizers are proposed for the linear and nonlinear case, based on sample complexity analysis
@@
Iwata et al. (2023) consider PDEs with polynomial governing equations. A single PINN is trained, which takes the query point and the coefficients of the PDE as input. Furthermore, Dirichlet boundary conditions are represented by a set of boundary points and their evaluations. The PINN is optimized with respect to the PINN-error, using self-generated data from randomly-sampled PDEs. However, to keep the number of coefficients small, the experiments are limited to two-dimensional PDEs with order 2 and use polynomials of degree 2.
@@
The aforementioned methods can be used to obtain initial model parameters that can serve as a starting point for fine-tuning. This fine-tuning is crucial for methods that learn a single initialization for all tasks, such as the MAML-based approaches. However, we also discussed methods that learn hyper-networks to predict initialization based on a given task specification. By using these methods without fine-tuning, we can construct zero-shot methods, that is, we can avoid training a new model when given a new PDE. In the next section, we consider neural processes, which are data-driven models that are particularly suitable for such fast adaptations.
@@ 4.3 Neural Processes
Hansen et al. (2023) consider the task of combining the data-driven NP with physical plausibility. They use an attentive NP as a solution prior, which provides an initial estimate of the solution through a multivariate Gaussian predictive distribution. To incorporate a soft physics-based conversation law constraint, they consider a linear constraint PDE of the form $\mathsf{L}u(t,x)=\int_{\mathcal{X}}u(t,x){\textnormal{d}}x=b(t)$, which can be approximated using discretization, which essentially translates obeying the soft physics constraint into solving a regression problem. By adopting a multivariate Gaussian solution prior, this prior can be conditioned on the constraint points to yield an updated distribution that is more physically plausible in closed form. The authors primarily consider the porous medium equation, which is used to model underground fluid flow, nonlinear heat transfer, and water desalination. This PDE can incorporate challenging solution constraints and discontinuities, which is why the solution’s physical plausibility is essential. The authors verify that the constraint correction step improves the solution towards the ground truth value.
@@ 5 Discussion on industrial applications
Across industry segments, a number of use cases gravitate around the prediction of wear in mechanical systems using meta-learning approaches (Chen et al., 2023; Kamranfar et al., 2021; Li et al., 2021c; Hu et al., 2021) as well as physics-informed losses (Hua et al., 2023).
@@
Anomaly detection Chen et al. (2023); Garouani et al. (2021a); Kim et al. (2022)
@@ 6 Conclusion
Additionally, there is a need for more research on uncertainty quantification, especially for industrial applications operating in data-scarce environments. Methods like neural processes (Section 4.3), which can be trained on extensive prior knowledge and fine-tuned with minimal contextual data, are particularly promising. These models not only adapt quickly but also provide valuable uncertainty estimates, which are crucial for informed decision-making in practical settings.
@@
Finally, it is inevitable that physics-based machine learning will follow the trend of foundation models Bommasani et al. (2021). This trend builds on the multi-task, meta, and contextual learning paradigms discussed in Section 4 by using large capacity transformer models that can learn from very large datasets, as seen for weather forecasting in Section 3.5. The goal of foundation models for physics-based ML is to learn a model across many settings, which can then be fine-tuned on new instances, potentially enabling data-efficient models through data-intensive pretraining (Subramanian et al., 2024; McCabe et al., 2023; Subramanian et al., 2024).
In summary, the integration of machine learning with classical methods, along with advancements in uncertainty quantification, will be essential for the continued development and application of physics-informed machine learning in both academic research and industrial applications.
@@ Appendix A Open-source Libraries and Datasets
Many more libraries exist. For neural differential equations, there is NeuroDiffGym/neurodiffeq [$554\bigstar$] by [Chen et al., 2020] in PyTorch. ‘Neural differential equations’ are essentially PINNs, but the authors cite the older literature. For PINNs, there is analysiscenter/pydens [$266\bigstar$] and Photon-AI-Research/NeuralSolvers [$118\bigstar$] in PyTorch, which implements a few extensions to PINN models. idrl-lab/idrlnet, also tackling PINNs in PyTorch, but can handle complex domain geometries. tensordiffeq/TensorDiffEq [$102\bigstar$] also implements PINNs, but using Tensorflow 2. In general, while these libraries appear impactful, they are typically older and less maintained as those discussed above.
For evaluation, pdebench/PDEBench [$545\bigstar$] by Takamoto et al. [2022] is a benchmark suite for PDE solvers to evaluate ML-based methods across a range of PDEs, parameters and boundary conditions.

### v2

@@ 4.2 Meta Learning
Whereas multi-task learning aims to exploit common knowledge while learning a given set of tasks in parallel, meta learning aims to optimize a learning procedure on the given tasks that can later be applied to novel tasks. Both paradigms can leverage a set of source tasks to learn better solutions on target tasks and are therefore closely related to transfer learning (Zhuang et al., 2021b). However, whereas multi-task learning optimizes source and target tasks jointly, meta learning considers the source tasks only during pre-training allowing for computationally efficient learning on the target tasks. In general, the result of meta-learning could be a complete algorithm. However, the methods discussed in the following are content with learning suitable initial parameters or other hyperparameters for given algorithms.
@@
Several works have applied these techniques for learning dynamics models (Lin et al., 2020; Qin et al., 2022). Meta-L (Lin et al., 2020) uses MAML for linear time-variant (LTV) systems. The model is optimized to achieve a small prediction error after updating the linear parameters of the dynamics with a single gradient step. During deployment, the fast adaptation makes it possible to adapt to changes of the dynamics online. MetaPDE (Qin et al., 2022) employs a similar approach to nonlinear dynamics. They compare MAML with a slightly different meta learning technique, LEAP (Flennerhag et al., 2019), for meta learning a PINN that can be quickly adapted to different PDE parameters. While LEAP was faster during meta-learning, MAML performed faster in deployment. NRPINN (Liu et al., 2022) is a similar approach that uses REPTILE (Nichol et al., 2018) on a modified PINN loss, which contains an additional term that can make use of labeled data, if available.
@@
Another MAML-based approach is used by iMODE (Li et al., 2023c), which adapts a latent input to a more general PINN that is shared among different PDE instances. A bi-level optimization problem is solved to optimize both, the common PINN network and the initialization of its latent input. Furthermore, PCA is applied to learn a compressed form of the latent input. When examples with labeled PDE parameters are available during training, iMode can learn a diffeomorphism between the labeled parameters and the compressed latent inputs, which can be used for system identification. DyAd (Wang et al., 2022d) directly feeds the parameters of the PDE as input to the model. The architecture is geared towards 2D images/velocity fields. It uses 3D convolutions and AdaPad, which predicts boundary conditions using a nonlinear transformation of the latent used for padding the images. As the PDE parameters are often unknown, DyAd also trains an encoder network that predicts the parameters from a history of states. The encoder is trained using weak supervision, e.g., by exploiting when a subset of the parameters are known.
@@
Instead of predicting latent inputs, hypernetworks can be used to predict the weights of the common PINN. For example, HyperPINN (de Avila Belbute-Peres et al., 2021) trains a hypernetwork that takes the parameters of a given class of PDEs (e.g., Burger’s equation) as input and outputs the parameters of a PINN. To reduce the number of the parameters of the PINN, Hyper-LR-PINNs (Cho et al., 2023b) restrict the weight matrices of the neural network to be low-rank, parameterized by the matrices of an SVD decomposition, that is, the weight matrix of layer $i$ is given by $\mathbf{W}_{i}=\mathbf{U}_{i}\bm{\Sigma}_{i}\mathbf{V}_{i}^{\top}$, with rectangular matrices $\mathbf{U}_{i}$ and $\mathbf{V}_{i}$, and a diagonal matrix $\bm{\Sigma}_{i}$. During offline pre-training, $\mathbf{U}$ and $\mathbf{V}$ are optimized directly along with the hypernetwork that predicts the elements of $\bm{\Sigma}$. During deployment, only the diagonal elements are optimized, using the hypernetwork for predicting their initial values. An alternative way to reduce the number of environment-specific parameters was proposed by Kirchmeyer et al. (2022). Their method, CoDA, learns shared nominal parameters of the dynamics model along with a shared linear projection matrix that maps environment-specific low-dimensional parameters to an offset for the nominal parameters. By penalizing the $\ell_{1}$ or $\ell_{2}$ norms of the projection matrix and the environment-specific parameters, CoDA aims to concentrate the resulting offset parameters around the nominal parameters for faster optimization during deployment.
@@
While the aforementioned methods were targeted towards a specific type of PDEs, some approaches aim to learn even more general networks. LEADS (Yin et al., 2021) learns a dynamics model, where the dynamics may also change over time. The model predicts the state derivatives and decomposes them into shared and environment-specific components. It is trained by minimizing the squared error between predicted derivatives and targets, with an additional convex regularizer penalizing the complexity of the environment-specific terms. Different regularizers are proposed for the linear and nonlinear cases, based on sample complexity analysis
@@
Iwata et al. (2023) consider PDEs with polynomial governing equations. A single PINN is trained, which takes the query point and the coefficients of the PDE as input. Furthermore, Dirichlet boundary conditions are represented by a set of boundary points and their evaluations. The PINN is optimized with respect to the PINN-error, using self-generated data from randomly sampled PDEs. However, to keep the number of coefficients small, the experiments are limited to two-dimensional PDEs of order 2 and use polynomials of degree 2.
@@
The aforementioned methods can be used to obtain initial model parameters that can serve as a starting point for fine-tuning. This fine-tuning is crucial for methods that learn a single initialization for all tasks, such as the MAML-based approaches. However, we also discussed methods that learn hyper-networks to predict initialization based on a given task specification (de Avila Belbute-Peres et al., 2021; Cho et al., 2023b; Kirchmeyer et al., 2022) . By using these methods without fine-tuning, we can construct zero-shot methods, that is, we can avoid training a new model when given a new PDE. In the next section, we consider neural processes, which are data-driven models that are particularly suitable for such fast adaptations.
@@ 4.3 Neural Processes
Hansen et al. (2023) consider the task of combining the data-driven NP with physical plausibility. They use an attentive NP as a solution prior, which provides an initial estimate of the solution through a multivariate Gaussian predictive distribution. To incorporate a soft physics-based conservation law constraint, they consider a linear constraint PDE of the form $\mathsf{L}\,u(t,x)=\int_{\mathcal{X}}u(t,x){\textnormal{d}}x=b(t)$, which can be approximated using discretization, which essentially translates obeying the soft physics constraint penalty into solving a regression problem subject to a weighting hyperparameter of the penalty. Note that this soft constraint becomes hard at the limit of the penalty weighting hyperparameter going to infinity. By adopting a multivariate Gaussian solution prior, this prior can be conditioned on the constraint points to yield an updated distribution that is more physically plausible in closed form. The authors primarily consider the porous medium equation, which is used to model underground fluid flow, nonlinear heat transfer, and water desalination. This PDE can incorporate challenging solution constraints and discontinuities, which is why the solution’s physical plausibility is essential. The authors verify that the constraint correction step improves the solution towards the ground truth value.
@@ 5 Discussion on industrial applications
Across industry segments, a number of use cases gravitate around the prediction of wear in mechanical systems using meta-learning approaches (Chen et al., 2023b; Kamranfar et al., 2021; Li et al., 2021c; Hu et al., 2021) as well as physics-informed losses (Hua et al., 2023).
@@
Anomaly detection Chen et al. (2023b); Garouani et al. (2021a); Kim et al. (2022)
@@ 6 Conclusion
Additionally, there is a need for more research on uncertainty quantification, especially for industrial applications operating in data-scarce environments. Methods like neural processes (section 4.3), which can be trained on extensive prior knowledge and fine-tuned with minimal contextual data, are particularly promising. These models not only adapt quickly but also provide valuable uncertainty estimates, which are crucial for informed decision-making in practical settings. This need is further emphasized in recent works such as Psaros et al. (2023), which provides a comprehensive overview of uncertainty quantification methods, metrics, and comparisons in scientific machine learning, and Mouli et al. (2024b), that explores the role of uncertainty quantification in improving out-of-domain generalization for PDE-based learning tasks. These contributions highlight the importance of probabilistic models like neural processes and Gaussian processes in addressing these challenges and represent a key area for future exploration.
@@
Finally, physics-based machine learning will inevitably follow the trend of foundation models (Bommasani et al., 2021). This trend builds on the multi-task, meta, and contextual learning paradigms discussed in section 4 by using large capacity transformer models that can learn from very large datasets, as seen for weather forecasting in section 3.5. The goal of foundation models for physics-based ML is to learn a model across many settings, which can then be fine-tuned on new instances, potentially enabling data-efficient models through data-intensive pretraining (Subramanian et al., 2024; McCabe et al., 2023; Herde et al., 2024).
This survey demonstrates that combining physics knowledge with machine learning has the power to reshape predictive modeling by building systems that are not only efficient and accurate but also generalizable and trustworthy. By leveraging centuries of scientific knowledge alongside the flexibility of modern machine learning, researchers can address problems that were previously deemed intractable.
As the field advances, the synergy between physics-based insights and data-driven methods will redefine scientific discovery and engineering design, unlocking solutions to challenges in domains as varied as climate modeling, materials science, and industrial process optimization. The future of this interdisciplinary field lies in striking the perfect balance between computational rigor, data efficiency, and practical applicability, ensuring its impact across both academic and industrial landscapes.
Acknowledgments
We wish to thank the anonymous reviewers for their feedback on this survey, and providing additional references to related work. This work was partly funded by Hessian.ai through the project ‘The Third Wave of Artificial Intelligence – 3AI’ by the Ministry for Science and Arts of the state of Hessen, by the grant “Einrichtung eines Labors des Deutschen Forschungszentrum für Künstliche Intelligenz (DFKI) an der Technischen Uni- versität Darmstadt”, and by the Hessian Ministry of Higher Education, Research, Science and the Arts (HMWK). O.W. acknowledges the financial support provided by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 492770117).
@@ Appendix A Open-source Libraries and Datasets
Many more libraries exist. For neural differential equations, there is NeuroDiffGym/neurodiffeq [$554\bigstar$] by [Chen et al., 2020] in PyTorch. However, somewhat confusingly, the ‘neural differential equations’ implemented in this library are essentially PINNs, but the authors cite the older literature. For PINNs, there is analysiscenter/pydens [$266\bigstar$] and Photon-AI-Research/NeuralSolvers [$118\bigstar$] in PyTorch, which implements a few extensions to PINN models. idrl-lab/idrlnet, also tackling PINNs in PyTorch, but can handle complex domain geometries. tensordiffeq/TensorDiffEq [$102\bigstar$] also implements PINNs, but using Tensorflow 2. In general, while these libraries appear impactful, they are typically older and less maintained as those discussed above.
For evaluation, pdebench/PDEBench [$545\bigstar$] by Takamoto et al. [2022] is a benchmark suite for PDE solvers to evaluate ML-based methods across a range of PDEs, parameters, and boundary conditions. Additionally, PDEArena [$241\bigstar$] by Gupta and Brandstetter [2022] provides an extensible platform for benchmarking PDE solvers, while APEBench [$59\bigstar$] focuses on adaptive physics-based machine learning techniques [Koehler et al., 2024]. nn-benchmark [$25\bigstar$] is designed for evaluating neural networks on scientific computing tasks [Otness et al., 2021], and The Well [$746\bigstar$] serves by Ohana et al. [2024] as a benchmark for geophysical modeling and machine learning in subsurface environments.